OpenStudy (anonymous):

(Using error theorem) For approximately what values of x can you replace sin x by x-(x^3/6) with an error magnitude no greater than 5x10^-4?

OpenStudy (anonymous):

@amistre64

OpenStudy (amistre64):

im guessing youve truncated the taylor series?

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

what does your error thrm tell us that helps us out then

OpenStudy (anonymous):

\[Rn(x)\le \frac{ f ^{n+1}(c)x ^{n+1} }{ (n+1)! }\]

OpenStudy (anonymous):

I was pretty sure that I got the right answer, but apparently not :/

OpenStudy (amistre64):

that seems to be an approximation for the 'at most' error. this is asking more for a specific value of x that will keep us in a valid range

OpenStudy (anonymous):

Ok. Well it said that there was a way to use the theorem. There's another theorem that they gave us, but we never use it in class, and my teacher says to use this one

OpenStudy (anonymous):

I got really close with my error formula

OpenStudy (anonymous):

But I guess I messed part of it up

OpenStudy (amistre64):

ok, its asking for a at most error perhaps we are looking for the value of c? does that seem valid?

OpenStudy (anonymous):

For c I used 1 because that is the most that the n+1th derivative will ever be

OpenStudy (anonymous):

My formula ended up being (x^(2n+3))/((2n+3)!)

OpenStudy (anonymous):

And if you plug in n=2, that spits out x^7/7!. The answer is apparently x^5/5!

OpenStudy (amistre64):

well, we did truncate at x^3, then next term in the series is the x^5 so the answer at least makes sense

OpenStudy (anonymous):

Oh shoot, I just realized that I am supposed to use n=4... Now that doesn't make any sense whatsoever

OpenStudy (anonymous):

I am extremely confused

OpenStudy (amistre64):

if we retain our terms 1 by 1, instead of only seeing the odd terms \[sin(x)=0+x-0x^2/2!-x^3/3!+0x^4/4!+x^5/5! -0x^6/6!-x^7/7!+...\] right?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

But I thought that I was able to generate a rule strictly for the nonzero terms. The sine maclaurin series ignores the zero terms

OpenStudy (amistre64):

it ignores them simply because it deals with (2n+1) .. it really isnt ignoring them is it? at any rate, the next term that was left out was the x^5 term, and the error thrm says that the |error| is no more than the first neglected term. if memory serves

OpenStudy (anonymous):

So my error formula has to account for the zero terms too?

OpenStudy (amistre64):

can you post your workings? or a screenshot of them?

OpenStudy (anonymous):

I mean all I did was\[R2(x)\le \frac{ 1(x ^{2(n+1)+1} }{ (2(n+1)+1)! }\]

OpenStudy (anonymous):

I forgot to close the parentheses on top

OpenStudy (anonymous):

But basically I just replaced n with n+1

OpenStudy (anonymous):

And I plugged it into my calculator with n=2, and it looked nothing like what it should've

OpenStudy (amistre64):

ok, n+1 is the next n value ... good and sin(x) or cos(x) is maxed at 1 good

OpenStudy (anonymous):

Ok. I can't figure out where I went wrong

OpenStudy (amistre64):

For approximately what values of x can you replace sin x by x-(x^3/6) with an error magnitude no greater than 5x10^-4 error thrm: the max error is no more than the first neglected term. the first neglected term is x^5/5! |x^5/5!| <= 5x10^-4 |x^5| <= 5!(5)(10^-4) |x| <= (5!(5)(10^-4))^(1/5) |x| <= .5697 this is how im reading the problem

OpenStudy (anonymous):

Well thats the answer

OpenStudy (anonymous):

I don't understand what you did though

OpenStudy (amistre64):

then im not sure why you did an n+1 and such

OpenStudy (amistre64):

what is the first neglected term?

OpenStudy (anonymous):

x^5/4!

OpenStudy (anonymous):

5!

OpenStudy (amistre64):

then, why not use it? when is x value of the first neglected term less than or equal to 5(10^-4) ...

OpenStudy (anonymous):

Why am I only concerned about the first neglected term?

OpenStudy (amistre64):

because the error thrm says so; and that is what you are spose to use. am i misstating the error thrm any?

OpenStudy (anonymous):

The theorem was that equation that I posted a while back

OpenStudy (amistre64):

yeah, at least the lagrange form of it.

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

That takes into account the infinite amount of terms that you cut off

OpenStudy (amistre64):

the sin is an alternating series isnt it ... (-1)^n

OpenStudy (anonymous):

Yes

OpenStudy (amistre64):

then im prolly remembering something about an alternating series which states what i stated. The error in an alternating series is no more than the first neglected term. \[sin(x)=\sum_{0}^{}\frac{(-1)^n}{(2n+1)!}x^{2n+1}\] \[sin(x)=x-\frac{x^3}{3!}+\sum_{2}^{}\frac{(-1)^n}{(2n+1)!}x^{2n+1}\]

OpenStudy (anonymous):

What do you mean that the error is no more than the first neglected term? We never learned anything like that

OpenStudy (amistre64):

you havent, but i have :) let me do some review, but notice that the Lagrange form also uses the first neglected term ....

OpenStudy (here_to_help15):

Ima try to explain without talking giberish i have solved this on another site a whole while back :)

OpenStudy (anonymous):

Ok thanks!!

OpenStudy (anonymous):

I know that it uses the first neglected term, but it accounts for all of the ones after that

OpenStudy (anonymous):

My question isn't why you are using the first neglected term. Its why you aren't accounting for all of the ones after that

OpenStudy (amistre64):

\[R_2(x)<=\left|(1)\frac{x^{2(2)+1}}{(2(2)+1)!}\right|\] we are using your langrange setup, and f^(n)(c)=1 regarless why is it you are not using this?

OpenStudy (amistre64):

by your own admission, the lagrange form accounts for all the other terms in this package right?

OpenStudy (amistre64):

the lagrange form utilizes the mean value thrm 

OpenStudy (anonymous):

I am using that. I was confused as to why you were only using x^5/5!

OpenStudy (amistre64):

i was using it out of some repressed memory lol

OpenStudy (anonymous):

And I am confused as to why my answer was wrong

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

you used x^7/7! instead of using the lagrange form correctly, you plugged in n=3 instead of n=2

OpenStudy (anonymous):

Is that really all I did.

OpenStudy (anonymous):

I spent over an hour trying to rectify this and that is all I did T_T

OpenStudy (anonymous):

I quit math

OpenStudy (amistre64):

\[\underbrace{\frac{x^{2(0)+1}}{(2(0)+1)!}-\frac{x^{2(1)+1}}{(2(1)+1)!}}_{=x-\frac{x^3}{6}}+\frac{x^{2(2)+1}}{(2(2)+1)!}-\]

OpenStudy (amistre64):

maybe you confused 6 as 6! and assumed the next term? i dunno ... its your psyche :)

OpenStudy (anonymous):

I just realized that you are completely right. I feel like a moron now XD

OpenStudy (anonymous):

Thank you so much for helping me!!